# On the Fundamental Limitations of Dual Static CVaR Decompositions in Markov Decision Processes

### v1

@@ A.1 Results of Section 3
Proof  First observe that, when fixing $\bm{\tilde{\xi}}$ and $\alpha$, the resulting risk-level at time $t$ for history $H\in\mathcal{H}_{T}$ is always given recursively by $Y_{t}(H|\bm{\tilde{\xi}},\alpha)=\alpha\prod_{\tau=1}^{t-1}\bm{\tilde{\xi}}(S_{\tau+1}|S_{\tau},A_{\tau},Y_{\tau})$ with initial condition $Y_{0}=\alpha$. To alleviate notation, we will mute these dependencies and simply write $Y_{t}$ when $H$, $\bm{\tilde{\xi}}$, and $\alpha$ are clear from context. Also note that we can exclude trajectories $H$ such that $P^{\tilde{\pi}_{h}^{\alpha}}(H)=0$ from our analysis because these do not induce any constraint in the definition of $\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ and they are always excluded from the computation of $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}$. As a result, in the rest of the proof we can always assume that we have $A_{t}=\tilde{\pi}(S_{t},Y_{T})$ when iterating over a history’s actions.
For a fixed $H\in\mathcal{H}_{T}$ and $\alpha$, we have by definition $\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)=\prod_{t=0}^{T-1}\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})$, which is nothing more than $Y_{T}/\alpha$ from our above remark. Because $\bm{\tilde{\xi}}(S_{t}|S_{t-1},Y_{t-1},A_{t-1})\in[0,1/Y_{t-1}]$ for all $t=1,\dots,T$, we also have that $Y_{t}\in[0,1]$, hence $Y_{T}\in[0,1]$. From the observation that $Y_{T}=\alpha\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)$, it follows that $\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)\in[0,1/\alpha]$ (i).
@@
=$\overset{(a)}{=}$ $\sum_{H\in\mathcal{H}_{T-1}}\sum_{S_{T}\in\mathcal{S}}\prod_{t=0}^{T-1}P(S_{t+1}|S_{t},A_{t})\,\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})$
=$\overset{(b)}{=}$ $\sum_{H\in\mathcal{H}_{T-1}}\prod_{t=0}^{T-2}P(S_{t+1}|S_{t},A_{t})\,\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})\sum_{S_{T}\in\mathcal{S}}P(S_{T}|S_{T-1},A_{T-1})\,\bm{\tilde{\xi}}(S_{T}|S_{T-1},Y_{T-1},A_{T-1})$
=$\overset{(c)}{=}$ $\sum_{H\in\mathcal{H}_{T-1}}\prod_{t=0}^{T-2}P(S_{t+1}|S_{t},A_{t})\,\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})$
For step (a) we used the observation that the set of histories of length $T$ is the set of histories at time $T-1$, upon which we concatenate the action $A_{T-1}$ selected by $\tilde{\pi}_{h}^{\alpha}$ and the possible next states $S_{T}$. In step (b) we exploited the fact that we can peel off the terms depending on $S_{T}$ from the product. Step (c) exploited the fact that $\bm{\tilde{\xi}}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a\right)$ for all $a=\tilde{\pi}(s,y)$. Repeating the above manipulation $T$ times yields
@@
yielding the desired property (ii). Combining properties (i) and (ii) proves $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$.
We now prove the stated equality by proving a slightly more general result. Let $\mathcal{H}_{T}^{s}$ denote the set of all histories $H\in\mathcal{H}_{T}$ beginning with $S_{0}=s$. We shall prove that the following result holds for all $s\in\mathcal{S}$ and $\alpha\in[0,1]$:
@@
Let $\alpha\in[0,1]$, $s\in\mathcal{S}$ and set $Y_{0}=\alpha$, we get
@@
where we used $A_{0}=\tilde{\pi}(s,Y_{0})$ and $y^{\prime}=Y_{0}\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})$ to alleviate notation. The result holds for all $s\in\mathcal{S}$ and $\alpha\in[0,1]$.
@@
=$\overset{(a)}{=}$ $\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,A_{0})\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})\left[\sum_{H^{\prime}\in\mathcal{H}_{t-1}^{s^{\prime}}}P^{\tilde{\pi}_{h}^{y^{\prime}}}(H^{\prime})\,\zeta_{y^{\prime}}(\bm{\tilde{\xi}})(H^{\prime})\left(\mathcal{R}(s,A_{0},s^{\prime})+\gamma\mathcal{R}_{0:t-1}^{H^{\prime}}\right)\right]$
=$\overset{(b)}{=}$ $\sum_{s^{\prime}\in\mathcal{S}}\sum_{H^{\prime}\in\mathcal{H}_{t-1}^{s^{\prime}}}P(s^{\prime}|s,A_{0})\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})\,P^{\tilde{\pi}_{h}^{y^{\prime}}}(H^{\prime})\,\zeta_{y^{\prime}}(\bm{\tilde{\xi}})(H^{\prime})\left(\mathcal{R}(s,A_{0},s^{\prime})+\gamma\mathcal{R}_{0:t-1}^{H^{\prime}}\right)$
=$\overset{(c)}{=}$ $\sum_{H\in\mathcal{H}_{t}^{s}}P^{\tilde{\pi}_{h}^{\alpha}}(H)\,\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)\,\mathcal{R}_{0:t}^{H},$
where we used $A_{0}=\tilde{\pi}(s,Y_{0})$ and $y^{\prime}=y_{0}\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})$ to alleviate notation. In step (a), we exploited the fact that $\sum_{H^{\prime}\in\mathcal{H}_{t-1}^{s^{\prime}}}P^{\tilde{\pi}_{h}^{y^{\prime}}}(H^{\prime})\,\zeta_{y^{\prime}}(\bm{\tilde{\xi}})(H^{\prime})=1$ to put the $\mathcal{R}(s,A_{0},s^{\prime})$ term inside the sum. Step (b) used the fact that the $P(s^{\prime}|s,A_{0})$ and $\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})$ terms do not rely on $H^{\prime}$ to move the summation. Finally, step (c) relied on the fact that all trajectories starting from $S_{0}=s$ can be expressed as the union over $s^{\prime}$ of all trajectories trajectories with $S_{1}=s^{\prime}$ and updating the necessary probability, perturbation, and reward terms accordingly. The result follows by setting $S=s_{0}$ and picking the desired $\alpha\in[0,1]$.
A.2 Results of Section 4
@@
Proof  We prove both results by direct construction.
($\bm{\implies}$) Assume $\xi$ is realizable by some $\bm{\tilde{\xi}}$. Define the assignment $\mathcal{Y}(H_{t})$ by recursively computing the risk-levels generated when following $\bm{\tilde{\xi}}$, that is $\mathcal{Y}(H_{0:t+1})=\mathcal{Y}(H_{0:t})\cdot\bm{\tilde{\xi}}(S_{t+1}|S_{t},\mathcal{Y}(H_{0:t}),A_{t})$ for all $t=0,\dots,T-1$ and initialized at $\mathcal{Y}(H_{0:0})=\alpha$. Because $\xi$ is realizable, it follows that $\mathcal{Y}(H)=\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)=\xi(H)$ for all histories $H\in\mathcal{H}_{T}$, establishing the risk propagation constraint. Now observe that by construction we have:
@@
By this observation and the definition of $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$, the state-level risk envelope and action selection constraints follow, proving the desired statement.
($\bm{\impliedby}$) Assume a consistent risk-level assignment $\mathcal{Y}$ exists. We construct $\bm{\tilde{\xi}}$ as follows: for any $(s,y,a)$ reached by a history $H$ with $P^{\tilde{\pi}_{h}^{\alpha}}(H)>0$, that is $s=S_{t},y=\mathcal{Y}(H_{0:t}),a=A_{t}$, define $\bm{\tilde{\xi}}(s^{\prime}|s,y,a)\coloneq\mathcal{Y}(H_{t}\cup(a,s^{\prime}))/y$. For all other $(s,y,a)$, define $\bm{\tilde{\xi}}(s^{\prime}|s,y,a)=1$ to trivially ensure it is part of $\Xi_{y}\left(P(\cdot|s,a)\right)$. By the state-level risk envelope and action selection constraints, it follows that $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}(\tilde{\pi})$. Moreover, for all histories we have $\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)=\prod_{t=0}^{T-1}\frac{\mathcal{Y}(H_{0:t+1})}{\mathcal{Y}(H_{0:t})}=\frac{\mathcal{Y}(H_{0:T})}{\mathcal{Y}(H_{0:0})}=\frac{\alpha\cdot\xi(H)}{\alpha}=\xi(H)$. Thus, $\xi$ is realizable.
@@
Proof  ($\bm{\implies}$) Assume $\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$. Let $\bm{\tilde{\xi}}^{\star}$ be a minimizer for $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s_{0},\alpha)$. From Proposition 5 we know we can compute $\xi^{\prime}\coloneq\zeta_{\alpha}(\bm{\tilde{\xi}}^{\star})$, where $\xi^{\prime}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ is a valid trajectory perturbation and $\sum_{H\in\mathcal{H}_{T}}P^{\tilde{\pi}_{h}^{\alpha}}(H)\,\xi^{\prime}(H)\,\mathcal{R}_{0:T}^{H}=\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}(s_{0},\alpha)=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$. Since we assumed $\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]=\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)$, this means $\xi^{\prime}$ must be an optimal trajectory perturbation. Because $\xi^{\prime}$ is realizable by construction, applying Lemma 8 gives the desired result.
($\bm{\impliedby}$) Let $\xi^{\star}$ be an optimal trajectory perturbation for which there exists a consistent risk assignment. By Lemma 8, we know $\xi^{\star}$ is realizable and hence there exists a $\bm{\tilde{\xi}}^{\prime}\in\bm{\tilde{\Xi}}(\tilde{\pi})$ such that $\zeta_{\alpha}(\bm{\tilde{\xi}}^{\prime})=\xi^{\star}$. From Proposition 5 and the definition of $\xi^{\star}$, we have $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\prime}}(s_{0},\alpha)=\sum_{H\in\mathcal{H}_{T}}P^{\tilde{\pi}_{h}^{\alpha}}(H)\,\xi^{\star}(H)\,\mathcal{R}_{0:T}^{H}=\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]$. Since $\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)=\min_{\bm{\tilde{\xi}}}\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s_{0},\alpha)$, it must be that $\bm{\upsilon}^{\tilde{\pi}}(s_{0},\alpha)\leq\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}^{\star}}(s_{0},\alpha)=\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]$. Combining this with the opposite inequality from Corollary 6 yields equality.
@@
Proof  ($\bm{\implies}$) Assume $\tilde{\pi}$ is uniformly optimal. By definition, for any $\alpha\in[0,1]$, we have $\operatorname{CVaR}_{\alpha}\left[Z^{\tilde{\pi}_{h}^{\alpha}}\right]=\max_{\pi_{h}}\operatorname{CVaR}_{\alpha}\left[Z^{\pi_{h}}\right]=\operatorname{CVaR}_{\alpha}\left[Z^{\pi_{h,\alpha}^{\star}}\right]$. Because we assumed there is only one optimal $\pi_{h}^{\star}(\alpha)$, it follows that the history distributions $P^{\tilde{\pi}_{h}^{\alpha}}$ and $P^{\pi_{h}^{\star}(\alpha)}$ must match. Because the history distributions only depend on policies in whether the action $A_{t}$ is selected, it follows that $\pi_{h}^{\alpha}$ and $\pi_{h}^{\star}(\alpha)$ always select the same action $A_{t}$ given $H_{0:t}$. Because of the feasibility of the dynamic program decomposition for $\pi_{h}(\alpha)$ (Corollary 9), this is equivalent to saying $\tilde{\pi}\big{(}S_{t},\mathcal{Y}_{\alpha}(H_{0:t})\big{)}=\pi_{h,\alpha}^{\star}(H_{0:t})$. As this must hold for all $\alpha\in[0,1]$, $\tilde{\pi}$ must satisfy the optimal-action-selection constraints.
($\bm{\impliedby}$) Assume $\tilde{\pi}$ satisfies the optimal-action-selection constraints. This means that for any given $\alpha\in[0,1]$ and any history $H_{0:t}$ reachable under the optimal policy $\pi_{h,\alpha}^{\star}$, the action chosen by $\tilde{\pi}_{h}^{\alpha}$ is the same as the action chosen by $\pi_{h,\alpha}^{\star}$. This ensures that the two policies are identical, $\tilde{\pi}_{h}^{\alpha}=\pi_{h,\alpha}^{\star}$. Consequently, their induced trajectory distributions are identical, $P^{\tilde{\pi}_{h}^{\alpha}}=P^{\pi_{h,\alpha}^{\star}}$. It follows directly that their CVaR evaluations must also be identical:

### v2

@@ B.1 Results of Section 3
Proof outline.
The proof proceeds in two steps. First, we establish that the mapping $\zeta_{\alpha}^{\bm{\tilde{\xi}}}$ produces valid history perturbations (i.e., within the risk envelope) by verifying that the cumulative product of state perturbations satisfies the required probability and boundedness constraints. Second, we prove the equality between the history-level summation and the recursive value function. This is achieved via backward induction, showing that the single-step updates aggregate exactly to the return weighted by the constructed history perturbations.
Proof.
We first prove $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ and then we prove $\sum_{H\in\mathcal{H}_{T}}P^{\tilde{\pi}_{h}^{\alpha}}(H)\,\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)\,\mathcal{R}_{0:T}^{H}=\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s_{0},\alpha)$.
Proof of $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$.
First observe that, when fixing $\bm{\tilde{\xi}}$ and $\alpha$, the resulting risk level at time $t$ for history $H\in\mathcal{H}_{T}$ is always given recursively by $Y_{t}(H|\bm{\tilde{\xi}},\alpha)=\alpha\prod_{\tau=1}^{t-1}\bm{\tilde{\xi}}(S_{\tau+1}|S_{\tau},A_{\tau},Y_{\tau})$ with initial condition $Y_{0}=\alpha$. To alleviate notation, we will mute these dependencies and simply write $Y_{t}$ when $H$, $\bm{\tilde{\xi}}$, and $\alpha$ are clear from context. Also note that we can exclude trajectories $H$ such that $P^{\tilde{\pi}_{h}^{\alpha}}(H)=0$ from our analysis because these do not induce any constraint in the definition of $\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$ and they are always excluded from the computation of $\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}$. As a result, in the rest of the proof we can always assume that we have $A_{t}=\tilde{\pi}(S_{t},Y_{T})$ when iterating over a history’s actions.
For a fixed $H\in\mathcal{H}_{T}$ and $\alpha$, we have by definition $\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)=\prod_{t=0}^{T-1}\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})$, which is nothing more than $Y_{T}/\alpha$ from our above remark. Because $\bm{\tilde{\xi}}(S_{t}|S_{t-1},Y_{t-1},A_{t-1})\in[0,1/Y_{t-1}]$ for all $t=1,\dots,T$, we also have that $Y_{t}=Y_{t-1}\bm{\tilde{\xi}}(S_{t}|S_{t-1},Y_{t-1},A_{t-1})\in(0,1]$. Hence we have in particular $Y_{T}\in(0,1]$. From the observation that $Y_{T}=\alpha\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)$, it follows that $\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)\in[0,1/\alpha]$ (i).
@@
$\overset{(a)}{=}$ $\sum_{H\in\mathcal{H}_{T-1}}\sum_{S_{T}\in\mathcal{S}}\prod_{t=0}^{T-1}P(S_{t+1}|S_{t},A_{t})\,\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})$
$\overset{(b)}{=}$ $\sum_{H\in\mathcal{H}_{T-1}}\prod_{t=0}^{T-2}P(S_{t+1}|S_{t},A_{t})\,\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})\sum_{S_{T}\in\mathcal{S}}P(S_{T}|S_{T-1},A_{T-1})\,\bm{\tilde{\xi}}(S_{T}|S_{T-1},Y_{T-1},A_{T-1})$
$\overset{(c)}{=}$ $\sum_{H\in\mathcal{H}_{T-1}}\prod_{t=0}^{T-2}P(S_{t+1}|S_{t},A_{t})\,\bm{\tilde{\xi}}(S_{t+1}|S_{t},Y_{t},A_{t})$
For step (a) we used the observation that the set of histories of length $T$ is the set of histories at time T1T-1, upon which we concatenate the action $A_{T-1}$ selected by $\tilde{\pi}_{h}^{\alpha}$ and the possible next states $S_{T}$. In step (b) we exploited the fact that we can peel off the terms depending on $S_{T}$ from the product. Step (c) exploited the fact that $\bm{\tilde{\xi}}(\cdot|s,y,a)\in\Xi_{y}\left(P(\cdot|s,a\right)$ for all $a=\tilde{\pi}(s,y)$. Repeating the above manipulation $T$ times yields
@@
, where we used the fact that $\bm{\tilde{\xi}}(\cdot|S_{0},Y_{0},A_{0})\in\Xi_{Y_{0}}\left(P(\cdot|S_{0},A_{0}\right)$ for the last equality, yielding the desired property (ii). Combining properties (i) and (ii) proves $\zeta_{\alpha}^{\bm{\tilde{\xi}}}\in\Xi_{\alpha}\left(P^{\tilde{\pi}_{h}^{\alpha}}\right)$.
Proof of $\sum_{H\in\mathcal{H}_{T}}P^{\tilde{\pi}_{h}^{\alpha}}(H)\,\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)\,\mathcal{R}_{0:T}^{H}=\bm{\upsilon}^{\tilde{\pi},\bm{\tilde{\xi}}}(s_{0},\alpha)$.
We now prove the stated equality by proving a slightly more general result. Let $\mathcal{H}_{T}^{s}$ denote the set of all histories $H\in\mathcal{H}_{T}$ beginning with $S_{0}=s$. We shall prove that the following result holds for all $s\in\mathcal{S}$ and $\alpha\in(0,1]$:
@@
Let $\alpha\in(0,1]$, $s\in\mathcal{S}$ and set $Y_{0}=\alpha$, we get
@@
where we used $A_{0}=\tilde{\pi}(s,Y_{0})$ and $y^{\prime}=Y_{0}\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})$ to alleviate notation. The result holds for all $s\in\mathcal{S}$ and $\alpha\in(0,1]$.
@@
$\overset{(a)}{=}$ $\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,A_{0})\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})\left[\sum_{H^{\prime}\in\mathcal{H}_{t-1}^{s^{\prime}}}P^{\tilde{\pi}_{h}^{y^{\prime}}}(H^{\prime})\,\zeta_{y^{\prime}}(\bm{\tilde{\xi}})(H^{\prime})\left(\mathcal{R}(s,A_{0},s^{\prime})+\gamma\mathcal{R}_{0:t-1}^{H^{\prime}}\right)\right]$
$\overset{(b)}{=}$ $\sum_{s^{\prime}\in\mathcal{S}}\sum_{H^{\prime}\in\mathcal{H}_{t-1}^{s^{\prime}}}P(s^{\prime}|s,A_{0})\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})\,P^{\tilde{\pi}_{h}^{y^{\prime}}}(H^{\prime})\,\zeta_{y^{\prime}}(\bm{\tilde{\xi}})(H^{\prime})\left(\mathcal{R}(s,A_{0},s^{\prime})+\gamma\mathcal{R}_{0:t-1}^{H^{\prime}}\right)$
$\overset{(c)}{=}$ $\sum_{H\in\mathcal{H}_{t}^{s}}P^{\tilde{\pi}_{h}^{\alpha}}(H)\,\zeta_{\alpha}^{\bm{\tilde{\xi}}}(H)\,\mathcal{R}_{0:t}^{H},$
where we used $A_{0}=\tilde{\pi}(s,Y_{0})$ and $y^{\prime}=y_{0}\,\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})$ to alleviate notation. In step (a), we exploited the fact that $\sum_{H^{\prime}\in\mathcal{H}_{t-1}^{s^{\prime}}}P^{\tilde{\pi}_{h}^{y^{\prime}}}(H^{\prime})\,\zeta_{y^{\prime}}(\bm{\tilde{\xi}})(H^{\prime})=1$ to put the $\mathcal{R}(s,A_{0},s^{\prime})$ term inside the sum. Step (b) used the fact that the $P(s^{\prime}|s,A_{0})$ and $\bm{\tilde{\xi}}(s^{\prime}|s,Y_{0},A_{0})$ terms do not rely on $H^{\prime}$ to move the summation. Finally, step (c) relied on the fact that all trajectories starting from $S_{0}=s$ can be expressed as the union over $s^{\prime}$ of all trajectories trajectories with $S_{1}=s^{\prime}$ and updating the necessary probability, perturbations, and reward terms accordingly. The result follows by setting $S=s_{0}$ and picking the desired $\alpha\in(0,1]$. ∎
B.2 Results of Section 4
@@
Proof outline.
[…]
Proof.
[…]
@@
By this observation and the definition of $\bm{\tilde{\xi}}\in\bm{\tilde{\Xi}}$, the state-level risk envelope and action selection constraints follow, proving the desired statement.
@@
Proof outline.
This proof relies on the equivalence established in Lemma 4 and the upper bound property from Corollary 1. For the forward direction, if the evaluation gap is zero, the optimal history perturbations must be realizable by some state-level perturbations, implying consistency. For the reverse direction, if consistency holds, the optimal history perturbations are realizable. This allows us to exactly match the dynamic program’s value between the true static CVaR and the realizable value, proving equality.
Proof.
[…]
Proof outline.
The result is proven by construction. We show that a Markovian policy can be trivially viewed as a risk-dependent policy where the action choice is independent of the risk level $y$. This renders the action-selection consistency constraints non-binding. Consequently, the standard CVaR decomposition theorem guarantees the existence of a valid risk assignment, ensuring no gap exists.
Proof.
We prove the statement by explicitly proving that $\tilde{\pi}=\pi$ is the desired policy. First observe that one can obtain a risk assignment $\mathcal{Y}$ by leveraging the CVaR Decomposition Theorem (Thm. 1 $T$ times, for instance by applying the dynamic programming operator in Equation 1. By construction, this risk assignment will satisfy the risk propagation and state-level risk envelope constraints. Considering $\tilde{\pi}=\pi$, the action-selection consistency constraints are also trivially satisfied, ensuring $\mathcal{Y}$ is a consistent risk assignment mapping from $\tilde{\pi}$ and $\pi$, hence $\tilde{\pi}$ does not have a CVaR evaluation gap as per Theorem 4.
B.3 Results of Section 5
[…]
Proof outline.
We prove the equivalence by linking the definitions of uniform optimality and action consistency. In the forward direction, uniform optimality implies the policy must match the optimal history-dependent policy for every $\alpha$, which in turn implies matching the actions prescribed by those optimal policies at the reachable risk-augmented states. In the reverse direction, satisfying the constraints ensures the policy mimics the optimal history-dependent policy for every $\alpha$, thereby guaranteeing the CVaR values match.
Proof.
[…]